OpenStudy (anonymous):

Graph the line y=3/2x-2

OpenStudy (anonymous):

this is of the standard form y=mx+c where m is the slope and c is the intercept, in this case m = 3/2 and c =-2. Check http://www.mathsisfun.com/equation_of_line.html

OpenStudy (anonymous):

\[y=\frac{3}{2 x-2} ? \]

OpenStudy (anonymous):

you can simplify it to (x-1) = 3/(2y) and then subsitute something like y'=1/y and plot y' vs x.